OpenStudy (anonymous):

How can ln(1/2) be rewritten as Log Rules as -log2?

OpenStudy (anonymous):

http://www.youtube.com/watch?v=4sF2IuP5R0w&feature=relmfu 43:29-33, the host states this.

OpenStudy (anonymous):

ln(1/2)=ln(1)-ln(2) and as ln(1)=0 you get -ln(2)